.O6X+160An investor decides at the beginning of the year to invest some of his cash in an account paying 8% annual interest, and to put the rest in a stock fund that ends up earning 15%over the course of the year. He puts $2000 more in the first account than in the stock fund, and at the end of the year he finds he has earned $1,310 in interest. How much money was invested at each of the two rates?

Answer by checkley71(8403)

(X+2000).08+.15X=1310 OR .O8X+160+.15X=1310 OR .23X=1310-160 OR .23X=1150 OR X=1150/23 OR X=5000 INVESTED @ 15% THE THE 8% INVESTMENT IS 5000+2000=7000
PROOF .08*7000+.15*5000=1310 OR 560+750=1310 OR 1310=1310